OpenStudy (anonymous):

Solve x = r2 - y2 for r

OpenStudy (aravindg):

solve for r2?

OpenStudy (anonymous):

solve for r

OpenStudy (aravindg):

is it \[\large r^2?\]

OpenStudy (aravindg):

did you mean \[x=r^2-y^2\]

OpenStudy (anonymous):

I'm sorry. Yes, x=sqrt of r squared minus sqrt of y squared, solve for r

OpenStudy (aravindg):

ok then add \(y^2\) on both sides what do you get ?

OpenStudy (anonymous):

first, how do you get the square beside the variable

OpenStudy (aravindg):

use equation button below

OpenStudy (anonymous):

uhhh nope...where

OpenStudy (anonymous):

ohhh, i see it now

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

x\[x ^{2}+y ^{2}=r ^{2}\]

OpenStudy (aravindg):

is it \[x^2?\]

OpenStudy (anonymous):

yes

OpenStudy (aravindg):

ok now you got equation for \(r^2\) .Now take square root on both sides to get r

OpenStudy (aravindg):

what will that be?

OpenStudy (anonymous):

Okay...I'm new to this site and I'm having a hard time with the "equation" button. I know the answer will be r = sqrt of x sqrd + y sqrd. My question...why does the square remove from the r and not the x and y?

OpenStudy (aravindg):

simply because square root does not distribute to each term

OpenStudy (aravindg):

\[\large \sqrt{a+b} \neq \sqrt{a} +\sqrt{b}\]

OpenStudy (aravindg):

i think that clears ur doubt

OpenStudy (anonymous):

Yes, it does. Thank you for ur patience. I'm going to play around with the equation buton. :-)

OpenStudy (aravindg):

lols all the best !! if you hav any problem about using the site ,fell free to ask me :)

OpenStudy (aravindg):

and @ericalashawn \[\huge \color{green}{\text{WELCOME TO OPENSTUDY!}}\]

OpenStudy (anonymous):

u are meant to say "GOTTA CATCH EM ALL!"

OpenStudy (aravindg):

lols @JayDS